with work please
In Problems 5 - 6, find the algebraic form for the following graphs of quadratics). 5: (1,5) (-2,-4) 6: TIL (1,3) (0,2)
Answer #1

soen 6 Given Vertex th, k) = (-2,-4) point on parabola = (15) standard form of parabola 9- 4 (1-1)^+k Putting volue of (h, k)Point (15) on pasobola hence it satisties the equ 5 = a (1+2)3-4 a/32 = 5+4 ga - 9 a=1 put Valve of a in equ o y = (x+2)2-4 y